# On world-sheet S-matrix of NSR string in static gauge

### v1

@@
As was shown in arXiv:1203.1054, expanding the Nambu action near the "long string" vacuum in the static gauge one finds that the one-loop $2\to 2$ scattering amplitude of the $D-2=24$ transverse 2d fluctuations $X^{i}$ is given by a pure phase expression consistent with underlying integrability. Similar computation in the Green-Schwarz superstring was carried out in arXiv:2404.09658. Here we consider the case of the NSR or spinning string starting with its manifestly 2d covariant action given by $D$ scalar multiplets coupled to 2d supergravity. Integrating out the zweibein and gravitino fields is not straightforward so that a Nambu-like version of the spinning string action involving only the scalar coordinates and their 2d fermionic partners was not given in the past. We show how the elimination of the auxiliary fields can be done in the static gauge after a particular choice of the superconformal gauges on the fermions while keeping the transverse fields $X^{i}$ and $\psi^{i}$ off-shell. The resulting one-loop S-matrix for $X^{i}$ turns out, as expected, to be the same as in the GS superstring case. We also discuss similar actions for the heterotic string and a $T\bar{T}$ deformation of a free 2d scalar multiplet.
@@ 1 Introduction
Starting with the Nambu action in flat $D$-dimensional space-time one may expand it near the "long string" vacuum $X^{a}=\xi^{a}$ ($a=0,1$). In the static gauge one gets an interacting action for the $D-2$ physical 2d massless scalars $X^{i}$
@@
One motivation to study the action of the spinning string in the static gauge is to compare it to a $T\bar{T}$ deformation of the free 2d scalar multiplet $(X^{i},\psi^{i})$ (cf. [11, 12, 13, 14]). We will comment on this in section 2.4 below.
@@ 2 Spinning string action expanded near long string vacuum
To compute the world-sheet S-matrix in the long string vacuum the first step is to expand the action (1.2) in the static gauge. This requires gauge fixing all local symmetries and integrating out the auxiliary fields to end up with an action for the physical $D-2$ components of the bosonic and fermionic fields $(X^{i},\psi^{i})$.
@@ 2.1 Gauge fixing
Let us first recall the notation we used in the spinning string action (1.2) (see also Appendix A). The world-sheet coordinates are $\xi^{a}$ ($a=0,1$), the induced metric is $g_{\mu\nu}=e_{\mu}^{a}e_{\nu}^{b}\eta_{ab},\ \eta_{ab}={\rm diag}(-1,1)$,   ($\mu,\nu=0,1$). The index $A=0,...,D-1$ labels the flat target space coordinates. $\psi^{A}$ are 2d Majorana fermions in real representation with $\bar{\psi}^{A}=(\psi^{A})^{T}\gamma^{0}$, and constant flat-space Dirac matrices ${\gamma}^{a}$ are related to ${\gamma}^{\mu}$ in (1.2) by $e^{a}_{\mu}$. We will denote the spinor components of $\psi$ as $\begin{pmatrix}\psi_{R}\\ \psi_{L}\end{pmatrix}$, with $\bar{\psi}=(\psi_{L},-\psi_{R})$ and similarly for $\chi_{\mu}$.
@@ 2.2 Solving for the auxiliary fields
$-\frac{1}{4}\bar{\psi}^{i}\gamma_{a}\partial_{b}\psi^{i}\bar{\psi}^{j}\gamma^{(a}\partial^{b)}\psi^{j}+\frac{1}{8}\bar{\psi}^{i}\gamma^{a}\partial_{{a}}\psi^{i}\bar{\psi}^{j}\gamma^{b}\partial_{b}\psi^{j}+....$
@@ 3.1 Heterotic string
$-\frac{1}{4}\bar{\Psi}^{{\mathcal{I}}}\gamma_{a}\partial_{b}\Psi^{{\mathcal{I}}}\bar{\Psi}^{{\mathcal{J}}}\gamma^{(a}\partial^{b)}\Psi^{{\mathcal{J}}}+\frac{1}{8}\bar{\Psi}^{{\mathcal{I}}}\gamma^{a}\partial_{{a}}\Psi^{{\mathcal{I}}}\bar{\Psi}^{{\mathcal{J}}}\gamma^{b}\partial_{b}\Psi^{{\mathcal{J}}}+...\ .$
@@ 3.2 GS superstring
$L_{1}=$ $-\sqrt{-\det\Big{[}\eta_{\mu\nu}+\partial_{\mu}X^{i}\partial_{\nu}X^{i}-2i\bar{\theta}^{I}\Gamma_{(\mu}\partial_{\nu)}\theta^{I}+...\Big{]}}$
$=$ $-1-\frac{1}{2}\partial_{\mu}X^{i}\partial^{\underline{\mu}}X^{i}+i\bar{\theta}^{I}\Gamma_{\mu}\partial^{\underline{\mu}}\theta^{I}$
$+\frac{1}{4}\partial_{\mu}X^{i}\partial_{\nu}X^{i}\partial^{\underline{\nu}}X^{j}\partial^{\underline{\mu}}X^{j}-i\bar{\theta}^{I}\Gamma_{(\underline{\mu}}\partial_{\nu)}\theta^{I}\partial^{\underline{\nu}}X^{j}\partial^{\underline{\mu}}X^{j}-\bar{\theta}^{I}\Gamma_{(\underline{\mu}}\partial_{\nu)}\theta^{I}\bar{\theta}^{J}\Gamma^{\underline{\mu}}\partial^{\underline{\nu}}\theta^{J}$
@@
$L_{2}=$ $-{i}\Big{[}\bar{\theta}^{1}\Gamma^{*}\Gamma^{\underline{\nu}}\partial_{\nu}\theta^{1}-\bar{\theta}^{2}\Gamma^{*}\Gamma^{\underline{\nu}}\partial_{\nu}\theta^{2}\Big{]}+{i}\epsilon^{\mu\nu}\partial_{\mu}X^{i}\Big{[}\bar{\theta}^{1}\Gamma^{i}\partial_{\nu}\theta^{1}-\bar{\theta}^{2}\Gamma^{i}\partial_{\nu}\theta^{2}\Big{]}+...\ ,$ (3.10)
where $\Gamma^{*}\equiv\Gamma^{0}\Gamma^{1}$. Defining the projectors $\mathcal{P}_{\pm}=\frac{1}{2}(1\pm\Gamma^{*})$ we can fix the $\kappa$-symmetry gauge as [4]
@@
Then the expansion does not contain cubic $\theta\theta X$ vertices and we get (omitting the explicit factors of projectors in (3.11)) [^10] Adding $L_{2}$ to $L_{1}$ gives gives the standard kinetic terms for $\theta^{1}$ and $\theta^{2}$ with extra $2P_{\pm}$ factors. The expression below is found after rescaling $\theta^{I}\rightarrow{1\over\sqrt{2}}\theta^{I}$ and omitting the explicit projector factors. We also use the flat-space indices as in (2.23).
@@
$-\frac{1}{4}\bar{\theta}^{I}\Gamma_{({a}}\partial_{b)}\theta^{I}\bar{\theta}^{J}\Gamma^{{a}}\partial^{{b}}\theta^{J}+\frac{1}{8}\bar{\theta}^{I}\Gamma^{{a}}\partial_{a}\theta^{I}\bar{\theta}^{J}\Gamma^{{b}}\partial_{b}\theta^{J}+...\ .$ (3.12)
Under the projection (3.11) $\theta^{I}$ have in total 16 real anticommuting components. Thus we get the same number of physical degrees of freedom and the same structure of the quartic interaction terms as in the spinning string case in (2.23). Note that the signs of the fermionic bilinear terms in (2.23) in (3.12) are opposite but they can be made the same by changing a representation of the gamma matrices.
@@ 3.3 ${T\bar{T}}$ deformation of free scalar multiplet
Given that the $(\partial X)^{4}$ terms in the expansion of the Nambu action in the static gauge can be viewed [27] as the $T\bar{T}$ deformation [28] of the free massless scalar theory one may wonder if a similar interpretation may apply also to the spinning string Lagrangian in (2.23). Starting with the free Lagrangian for $D-2$ scalar multiplets on a flat 2d background
@@
Then the leading order of the $T\bar{T}$ deformation (${\partial\mathcal{L}\over\partial\lambda}={1\over 2}\det T_{ab}$) is proportional to
$\mathcal{L}_{1}={1\over 2}\det T_{ab}=\frac{1}{4}(T^{a}_{a}T^{b}_{b}-T^{a}_{b}T^{b}_{a})=$ $\ \ \frac{1}{4}\partial_{a}X^{i}\partial_{b}X^{i}\partial^{a}X^{j}\partial^{b}X^{j}-\frac{1}{8}\partial_{a}X^{i}\partial^{a}X^{i}\partial_{b}X^{j}\partial^{b}X^{j}$ (3.15)
@@
The sum $\mathcal{L}=\mathcal{L}_{0}+\mathcal{L}_{1}$ is thus equivalent to (2.23). The corresponding action is invariant under the modified global supersymmetry relating $X^{i}$ and $\psi^{i}$ as discussed in Appendix B.
@@ 4 One-loop scattering amplitude
$(p_{1},i)$$(p_{2},j)$$(p_{3},k)$$(p_{4},l)$$s$-channel$(p,a)$$(q,b)$
$(p_{1},i)$$(p_{2},j)$$(p_{3},k)$$(p_{4},l)$$t$-channel$(p,a)$$(q,b)$
$(p_{1},i)$$(p_{2},j)$$(p_{4},k)$$(p_{3},l)$$u$-channel$(p,a)$$(q,b)$
@@
The massless 2d kinematics implies $stu=0$ so that choosing $t=0,\ u=-s$ we get $A^{(0)}=C^{(0)}=0$, $B^{(0)}=-\frac{1}{2}s^{2}$.
@@
$A_{b}^{(1)}=-C_{b}^{(1)}=-\frac{{\hat{D}}-24}{192\pi}s^{3}\ ,\ \ \ \ \ \ \qquad B_{b}^{(1)}=\frac{i}{16}s^{3},$ (4.4)
@@
[…]
@@
where $\Delta$ was defined in (4.8). As a result, the fermionic loop contribution to the $s$-channel $A$-amplitude can be written as
@@
The expected agreement between the spinning string and GS superstring amplitudes follows from the close similarity of the corresponding 4-vertices in the Lagrangians in (2.23) and (3.12). In the GS case $\theta^{I}$ subject to (3.11) have 16 independent components. In the spinning string case $\psi^{i}$ has $n_{f}\,\hat{D}=2(D-2)$ components which for $D=10$ gives the same factor of 16 in the fermionic loop. [^13] It is interesting to note that the coefficient of the UV divergent part of (4.18) (corresponding to the total derivative curvature scalar $\sqrt{-g}R^{(2)}$ term evaluated on the induced metric) given for $n_{f}=2$ by ${3\over 2}\hat{D}-6={3\over 2}(D-6)$ is the same as found in the partition function of the spinning string on a general curved 2d background [30].
@@ Acknowledgements
We thank F. Seibold for useful discussions and sharing a Mathematica code for computing one-loop amplitudes in the GS string case. AAT is grateful to R. Metsaev for important discussions and suggestions. ZW would like to thank H. Jiang, K. Mkrtchyan and D. Zhong for many discussions. We also thank F. Seibold and R. Metsaev for comments on the draft. This work was supported by the STFC grant ST/T000791/1.
@@ Appendix A Conventions and useful relations
$\bar{\psi}^{a}\gamma^{1}\partial_{b}\psi_{a}=\psi_{L}^{a}\partial_{b}\psi_{a,L}-\psi_{R}^{a}\partial_{b}\psi_{a,R}=-\frac{1}{2}\psi_{L}^{+}\partial_{b}\psi^{-}_{L}-\frac{1}{2}\psi_{L}^{-}\partial_{b}\psi^{+}_{L}+\frac{1}{2}\psi_{R}^{+}\partial_{b}\psi^{-}_{R}+\frac{1}{2}\psi_{R}^{-}\partial_{b}\psi^{+}_{R}\ ,$ (A.6)
$\bar{\psi}^{a}\gamma^{b}\partial_{b}\psi_{a}=-2(\psi_{L}^{a}\partial_{-}\psi_{a,L}+\psi_{R}^{a}\partial_{+}\psi_{a,R})=\psi_{L}^{+}\partial_{-}\psi^{-}_{L}+\psi_{L}^{-}\partial_{-}\psi^{+}_{L}+\psi_{R}^{+}\partial_{+}\psi^{-}_{R}+\psi_{R}^{-}\partial_{+}\psi^{+}_{R}\ .$ (A.7)

### v4

@@
As was shown in arXiv:1203.1054, expanding the Nambu action near the "long string" vacuum in the static gauge one finds that the one-loop $2\to 2$ scattering amplitude of the D2=24D-2=24 transverse 2d fluctuations $X^{i}$ is given by a pure phase expression consistent with underlying integrability. Similar computation in the Green-Schwarz superstring was carried out in arXiv:2404.09658. Here we consider the case of the NSR or spinning string starting with its manifestly 2d covariant action given by $D$ scalar multiplets coupled to 2d supergravity. Integrating out the zweibein and gravitino fields is not straightforward so that a Nambu-like version of the spinning string action involving only the scalar coordinates and their 2d fermionic partners was not given in the past. We show how the elimination of the auxiliary fields can be done in the static gauge after a particular choice of the superconformal gauges on the fermions while keeping the transverse fields $X^{i}$ and $\psi^{i}$ off-shell. The resulting one-loop S-matrix for $X^{i}$ turns out, as expected, to be the same as in the GS superstring case. We also discuss similar actions for the heterotic string and a $T\bar{T}$ deformation of a free 2d scalar multiplet.
@@ 1 Introduction
Starting with the Nambu action in flat $D$-dimensional space-time one may expand it near the "long string" vacuum $X^{a}=\xi^{a}$ ($a=0,1$). In the static gauge one gets an interacting action for the D2D-2 physical 2d massless scalars $X^{i}$
@@
One motivation to study the action of the spinning string in the static gauge is to compare it to a $T\bar{T}$ deformation of the free 2d scalar multiplet $(X^{i},\psi^{i})$ (cf. [11, 12, 13, 14, 15, 16, 17]). We will comment on this in section 2.4 below.
@@ 2 Spinning string action expanded near long string vacuum
To compute the world-sheet S-matrix in the long string vacuum the first step is to expand the action (1.2) in the static gauge. This requires gauge fixing all local symmetries and integrating out the auxiliary fields to end up with an action for the physical D2D-2 components of the bosonic and fermionic fields $(X^{i},\psi^{i})$.
@@ 2.1 Gauge fixing
Let us first recall the notation we used in the spinning string action (1.2) (see also Appendix A). The world-sheet coordinates are $\xi^{a}$ ($a=0,1$), the induced metric is $g_{\mu\nu}=e_{\mu}^{a}e_{\nu}^{b}\eta_{ab},\ \eta_{ab}={\rm diag}(-1,1)$,   ($\mu,\nu=0,1$). The index A=0,,D1A=0,...,D-1 labels the flat target space coordinates. $\psi^{A}$ are 2d Majorana fermions in real representation with $\bar{\psi}^{A}=(\psi^{A})^{T}\gamma^{0}$, and constant flat-space Dirac matrices ${\gamma}^{a}$ are related to ${\gamma}^{\mu}$ in (1.2) by $e^{a}_{\mu}$. We will denote the spinor components of $\psi$ as $\begin{pmatrix}\psi_{R}\\ \psi_{L}\end{pmatrix}$, with $\bar{\psi}=(\psi_{L},-\psi_{R})$ and similarly for $\chi_{\mu}$.
@@ 2.2 Solving for the auxiliary fields
$-\frac{1}{4}\bar{\psi}^{i}\gamma_{a}\partial_{b}\psi^{i}\bar{\psi}^{j}\gamma^{(a}\partial^{b)}\psi^{j}+\frac{1}{8}\bar{\psi}^{i}\gamma^{a}\partial_{{a}}\psi^{i}\bar{\psi}^{j}\gamma^{b}\partial_{b}\psi^{j}+.$.
@@ 3.1 Heterotic string
$-\frac{1}{4}\bar{\Psi}^{{\mathcal{I}}}\gamma_{a}\partial_{b}\Psi^{{\mathcal{I}}}\bar{\Psi}^{{\mathcal{J}}}\gamma^{(a}\partial^{b)}\Psi^{{\mathcal{J}}}+\frac{1}{8}\bar{\Psi}^{{\mathcal{I}}}\gamma^{a}\partial_{{a}}\Psi^{{\mathcal{I}}}\bar{\Psi}^{{\mathcal{J}}}\gamma^{b}\partial_{b}\Psi^{{\mathcal{J}}}+.\$.
@@ 3.2 GS superstring
$L_{1}=$ $-\sqrt{-\det\Big{[}\eta_{\mu\nu}+\partial_{\mu}X^{i}\partial_{\nu}X^{i}-2i\bar{\theta}^{I}\Gamma_{(\underline{\mu}}\partial_{\nu)}\theta^{I}-\bar{\theta}^{I}\Gamma_{\underline{\mu}}\partial_{\nu}\theta^{I}\bar{\theta}^{J}\Gamma^{\underline{\mu}}\partial^{\underline{\nu}}\theta^{J}+...\Big{]}}$
$=$ $-1-\frac{1}{2}\partial_{\mu}X^{i}\partial^{\underline{\mu}}X^{i}+i\bar{\theta}^{I}\Gamma^{\mu}\partial_{\underline{\mu}}\theta^{I}$
$+\frac{1}{4}\partial_{\mu}X^{i}\partial_{\nu}X^{i}\partial^{\underline{\mu}}X^{j}\partial^{\underline{\nu}}X^{j}-i\bar{\theta}^{I}\Gamma_{\underline{\mu}}\partial_{\nu}\theta^{I}\partial^{\underline{\mu}}X^{j}\partial^{\underline{\nu}}X^{j}-\frac{1}{2}\bar{\theta}^{I}\Gamma_{\underline{\nu}}\partial_{\mu}\theta^{I}\bar{\theta}^{J}\Gamma^{\underline{\mu}}\partial^{\underline{\nu}}\theta^{J}$
@@
$L_{2}=$ $-{i}\Big{[}\bar{\theta}^{1}\Gamma^{*}\Gamma^{\underline{\nu}}\partial_{\nu}\theta^{1}-\bar{\theta}^{2}\Gamma^{*}\Gamma^{\underline{\nu}}\partial_{\nu}\theta^{2}\Big{]}+\epsilon^{\mu\nu}\bar{\theta}^{1}\Gamma^{\underline{\rho}}\partial_{\mu}\theta^{1}\bar{\theta}^{2}\Gamma_{\underline{\rho}}\partial_{\nu}\theta^{2}+$...\ , (3.10)
Defining the projectors $\mathcal{P}_{\pm}=\frac{1}{2}(1\pm\Gamma^{*})$ we can fix the $\kappa$-symmetry gauge as [4]
@@
Then the expansion does not contain cubic $\theta\theta X$ vertices and we get (omitting the explicit factors of projectors in (3.11)) [^10] Adding $L_{2}$ to $L_{1}$ gives gives the standard kinetic terms for $\theta^{1}$ and $\theta^{2}$ with extra $2P_{\pm}$ factors. The expression below is found after rescaling $\theta^{I}\rightarrow{1\over\sqrt{2}}\theta^{I}$ and omitting the explicit projector factors. We also use the flat-space indices as in (2.23). We thank M. Beccaria and R. Roiban for correcting the coefficients of the $\theta^{4}$ terms.
@@
$-\frac{1}{4}\bar{\theta}^{I}\Gamma_{{a}}\partial_{b}\theta^{I}\bar{\theta}^{J}\Gamma^{{b}}\partial^{{a}}\theta^{J}+\frac{1}{4}\bar{\theta}^{I}\Gamma^{{a}}\partial_{a}\theta^{I}\bar{\theta}^{J}\Gamma^{{b}}\partial_{b}\theta^{J}+.$..\ . (3.12)
Under the projection (3.11) $\theta^{I}$ have in total 16 real anticommuting components. Thus we get the same number of physical degrees of freedom and the same structure of the quartic interaction terms as in the spinning string case in (2.23). [^11] Note that the signs of the fermionic bilinear terms in (2.23) in (3.12) are opposite but they can be made the same by changing a representation of the gamma matrices.
Note that (3.12) contains only $\Gamma_{a}$ matrices with 2d indices and thus one can chose their representation so that to convert $\theta^{I}$ into a collection of eight 2d spinors (in particular, $\bar{\theta}^{I}\Gamma^{a}\partial_{b}\theta^{I}$ written explicitly in components becomes the same as $\bar{\psi}^{i}\gamma^{a}\partial_{b}\psi^{i}$). Redefining $\psi^{i}$ in (2.23) as
$\psi^{i}\rightarrow\psi^{i}-\frac{i}{16}\big{[}\bar{\psi}^{j}\gamma^{a}\partial_{a}\psi^{j}+(\bar{\psi}^{j}\gamma^{*}\gamma^{a}\partial_{a}\psi^{j})\gamma^{*}\big{]}\,\psi^{i}+\dots,$ (3.13)
one can establish direct equivalence between the NSR action (2.23) and the GS one in (3.12). This shows that the correspondence between the two actions familiar in the light-cone gauge holds also in the static gauge.
@@ 3.3 $T\bar{T}$} deformation of free scalar multiplet
Given that the $(\partial X)^{4}$ terms in the expansion of the Nambu action in the static gauge can be viewed [30] as the $T\bar{T}$ deformation [31] of the free massless scalar theory one may wonder if a similar interpretation may apply also to the spinning string Lagrangian in (2.23). Starting with the free Lagrangian for D2D-2 scalar multiplets on a flat 2d background
@@
Then the leading order of the $T\bar{T}$ deformation (${\partial\mathcal{L}\over\partial\lambda}=-{1\over 2}\det T_{ab}$) is proportional to
$\mathcal{L}_{1}=-{1\over 2}\det T_{ab}=-\frac{1}{4}(T^{a}_{a}T^{b}_{b}-T^{a}_{b}T^{b}_{a})=$ $\ \ \frac{1}{4}\partial_{a}X^{i}\partial_{b}X^{i}\partial^{a}X^{j}\partial^{b}X^{j}-\frac{1}{8}\partial_{a}X^{i}\partial^{a}X^{i}\partial_{b}X^{j}\partial^{b}X^{j}$ (3.16)
@@
The sum $\mathcal{L}=\mathcal{L}_{0}+\mathcal{L}_{1}$ is thus equivalent to (2.23). The corresponding action is invariant under the modified global supersymmetry relating $X^{i}$ and $\psi^{i}$ as discussed in Appendix B. [^13] Let us note that the expansion of the GS action in (3.12) can also be reproduced as a similar $T\bar{T}$ deformation as in (3.16) if $T_{ab}$ is chosen as the canonical stress tensor corresponding to the quadratic part of (3.12).
@@ 4 One-loop scattering amplitude
$(p_{1},i)$$(p_{2},j)$$(p_{3},k)$$(p_{4},l)$$s$-channel(p,a)(p,a)$(q,b)$
$(p_{1},i)$$(p_{2},j)$$(p_{3},k)$$(p_{4},l)$$t$-channel(p,a)(p,a)$(q,b)$
$(p_{1},i)$$(p_{2},j)$$(p_{4},k)$$(p_{3},l)$$u$-channel(p,a)(p,a)$(q,b)$
@@
The massless 2d kinematics implies $stu=0$ so that choosing t=0,u=st=0,\ u=-s we get $A^{(0)}=C^{(0)}=0$, $B^{(0)}=-\frac{1}{2}s^{2}$.
@@
$A_{b}^{(1)}=-C_{b}^{(1)}=-\frac{{\hat{D}}-24}{192\pi}s^{3}\ ,\ \ \ \ \ \ \qquad B_{b}^{(1)}=\frac{i}{16}s^{3},\ \$(4.4)
@@
$A^{(1)}_{f,s}=\frac{1}{2i}\int\frac{d^{d}p}{(2\pi)^{d}}\frac{N_{A,s}}{(p^{2}-i\varepsilon)(q^{2}-i\varepsilon)}=\frac{1}{2i}\int^{1}_{0}dx\int\frac{d^{d}l}{(4\pi)^{d}}\frac{N_{A,s}}{(l^{2}+\Delta)^{2}},$ (4.6)
@@
where $\Delta$ was defined in (4.8). As a result, the fermionic loop contribution to the $s$-channel AA-amplitude can be written as
@@
The expected agreement between the spinning string and GS superstring amplitudes follows from the close similarity of the corresponding 4-vertices in the Lagrangians in (2.23) and (3.12). In the GS case $\theta^{I}$ subject to (3.11) have 16 independent components. In the spinning string case $\psi^{i}$ has $n_{f}\,\hat{D}=2(D-2)$ components which for $D=10$ gives the same factor of 16 in the fermionic loop.
@@ Acknowledgements
We thank F. Seibold for useful discussions and sharing a Mathematica code for computing one-loop amplitudes in the GS string case. AAT is grateful to R. Metsaev for important discussions and suggestions. ZW would like to thank H. Jiang, K. Mkrtchyan and D. Zhong for many discussions. We also thank R. Metsaev and F. Seibold for comments on the draft. This work was supported by the STFC grant ST/T000791/1.
@@ Appendix A Conventions and useful relations
$\bar{\psi}^{a}\gamma^{1}\partial_{b}\psi_{a}=\psi_{L}^{a}\partial_{b}\psi_{a,L}-\psi_{R}^{a}\partial_{b}\psi_{a,R}=-\frac{1}{2}\psi_{L}^{+}\partial_{b}\psi^{-}_{L}-\frac{1}{2}\psi_{L}^{-}\partial_{b}\psi^{+}_{L}+\frac{1}{2}\psi_{R}^{+}\partial_{b}\psi^{-}_{R}+\frac{1}{2}\psi_{R}^{-}\partial_{b}\psi^{+}_{R}\$, (A.6)
$\bar{\psi}^{a}\gamma^{b}\partial_{b}\psi_{a}=-2(\psi_{L}^{a}\partial_{-}\psi_{a,L}+\psi_{R}^{a}\partial_{+}\psi_{a,R})=\psi_{L}^{+}\partial_{-}\psi^{-}_{L}+\psi_{L}^{-}\partial_{-}\psi^{+}_{L}+\psi_{R}^{+}\partial_{+}\psi^{-}_{R}+\psi_{R}^{-}\partial_{+}\psi^{+}_{R}\$. (A.7)